# LongCodeBench: Evaluating Coding LLMs at 1M Context Windows

### v1

@@ B.1 Prompt to identify issues convertible into questions
You are an expert software engineer and teacher. You
are given a closed GitHub issue from a public repository and the comments on
the issue. Read the issue and assess whether it is simple enough that its
resolution boils down to clarifying a concept about the codebase. If so,
reexpress it as a multiple choice question with a correct answer and three
wrong answers.
Issue whose resolution required a pull request or a code change are examples
of issues that are not simple enough for this task.
Moreover, the question should be about both the underlying concept and the
codebase, not a general question about programming, software engineering, or
the language being used.
Examples of issues that are simple enough to be expressed as multiple choice
questions:
- Clarifying the behavior of a function, method, or module.
- Understanding the purpose of a variable or class.
- Explaining the flow of control in a code snippet.
- Explaining why a certain error occurs in a code snippet due to a subtle, but
wanted, behavior of the library being used.
Examples of issues that are not simple enough to be expressed as multiple
choice questions:
- Fixing a bug in the code.
- Implementing a new feature.
- Suggestions on what to pair with a certain library.
- Installation issues.
- Questions about the project’s roadmap.
- Observations relating to software engineering practices.
Please provide reasoning to justify your decision. Be conservative, and only
consider issues as simple enough if it is clear beyond doubt that they are.
Issue {issue["title"]}: {issue["body"]}
Comments: {issue["comments"]}
@@ B.2 Prompt to filter questions requiring repository-specific knowledge
You are provided with a question about the repository {repo}. Your task is not to
answer the question directly but to evaluate whether an accurate answer requires a
detailed, up-to-date understanding of the repository, or if it can be answered
accurately using only your pre-existing knowledge (which may include exposure to the
repository during training) and general programming knowledge.
Please explain your reasoning in detail. At the end of your response, on a new line,
output only ’Yes’ if you believe the question requires repository-specific knowledge
to answer accurately, or ’No’ if you believe it can be answered correctly without
directly consulting the repository.
Here is the question:
{question and answers}
@@ B.3 Prompt for LongCodeQA task
You are a coding expert. Your task is to analyze a GitHub repository and then answer
one question about it.
Repository:
{repo_text}
Question:
{question}
Please analyze the repository text, reason through the question, and then choose
among A, B, C, D answers the correct one.
Provide your final answer in the following format:
Final Answer: <LETTER>
@@ B.4 Prompt for LongSWE-Bench task
You will be provided with a partial code base and an issue statement explaining a
problem to resolve.
<issue>
{issue_body}
</issue>
<code>
{repo_body}
</code>
Here is an example of a patch file. It consists of changes to the code base. It
specifies the file names, the line numbers of each change, and the removed and added
lines. A single patch file can contain changes to multiple files.
<patch>
--- a/file.py
+++ b/file.py
@@ -1,27 +1,35 @@
def euclidean(a, b):
- while b:
- a, b = b, a % b
- return a
+ if b == 0:
+ return a
+ return euclidean(b, a % b)
[...truncated for brevity...]
</patch>
I need you to solve the provided issue by generating a single patch file that I can
apply directly to this repository using git apply. Please respond with a single patch
file in the format shown above.
Respond below:
@@ Appendix C Inference time and cost
| Claude 3.5 Sonnet | Anthropic | 100USD | 140USD | |

### v2

@@ B.1 Prompt to identify issues convertible into questions
You are an expert software engineer and teacher. You are given a closed GitHub issue from a public repository and the comments on the issue. Read the issue and assess whether it is simple enough that its resolution boils down to clarifying a concept about the codebase. If so, reexpress it as a multiple choice question with a correct answer and three wrong answers. Issue whose resolution required a pull request or a code change are examples of issues that are not simple enough for this task. Moreover, the question should be about both the underlying concept and the codebase, not a general question about programming, software engineering, or the language being used. Examples of issues that are simple enough to be expressed as multiple choice questions: - Clarifying the behavior of a function, method, or module. - Understanding the purpose of a variable or class. - Explaining the flow of control in a code snippet. - Explaining why a certain error occurs in a code snippet due to a subtle, but wanted, behavior of the library being used. Examples of issues that are not simple enough to be expressed as multiple choice questions: - Fixing a bug in the code. - Implementing a new feature. - Suggestions on what to pair with a certain library. - Installation issues. - Questions about the project’s roadmap. - Observations relating to software engineering practices. Please provide reasoning to justify your decision. Be conservative, and only consider issues as simple enough if it is clear beyond doubt that they are. Issue {issue["title"]}: {issue["body"]} Comments: {issue["comments"]}
@@ B.2 Prompt to filter questions requiring repository-specific knowledge
You are provided with a question about the repository {repo}. Your task is not to answer the question directly but to evaluate whether an accurate answer requires a detailed, up-to-date understanding of the repository, or if it can be answered accurately using only your pre-existing knowledge (which may include exposure to the repository during training) and general programming knowledge. Please explain your reasoning in detail. At the end of your response, on a new line, output only ’Yes’ if you believe the question requires repository-specific knowledge to answer accurately, or ’No’ if you believe it can be answered correctly without directly consulting the repository. Here is the question: {question and answers}
@@ B.3 Prompt for LongCodeQA task
You are a coding expert. Your task is to analyze a GitHub repository and then answer one question about it. Repository: {repo_text} Question: {question} Please analyze the repository text, reason through the question, and then choose among A, B, C, D answers the correct one. Provide your final answer in the following format: Final Answer: <LETTER>
@@ B.4 Prompt for LongSWE-Bench task
You will be provided with a partial code base and an issue statement explaining a problem to resolve. <issue> {issue_body} </issue> <code> {repo_body} </code> Here is an example of a patch file. It consists of changes to the code base. It specifies the file names, the line numbers of each change, and the removed and added lines. A single patch file can contain changes to multiple files. <patch> --- a/file.py +++ b/file.py @@ -1,27 +1,35 @@ def euclidean(a, b): - while b: - a, b = b, a % b - return a + if b == 0: + return a + return euclidean(b, a % b) [...truncated for brevity...] </patch> I need you to solve the provided issue by generating a single patch file that I can apply directly to this repository using git apply. Please respond with a single patch file in the format shown above. Respond below:
@@ Appendix C Inference time and cost
Claude 3.5 Sonnet | Anthropic | 100USD | 140USD |